A young couple, Eve and Steven, want to divide their main household chores (budgeting, cooking, dishwashing,...

  1. A young couple, Eve and Steven, want to divide their main household chores (budgeting, cooking, dishwashing, and laundering) between them so that each has two tasks but the total time they spend on household duties is kept to a minimum. Their efficiencies on these tasks differ, where the time each would need to perform the task is given by the following table:

Budgeting (1)

Cooking(2)

Dishwashing(3)

Laundry(4)

Eve

4.5 hrs

7.8 hrs

3.6 hrs

2.9 hrs

Steven

4.9 hrs

7.2 hrs

4.3 hrs

3.1 hrs

Formulate this problem as a linear programming problem and solve it using MS Excel Solver.
Answer #1

Formulation:

Let Yjk be the binary integer such that the j-th person is appointed to the k-th job for j=1,2 and k=1,2,3,4

Minimize Z = total time = 4.5 Y11 + 7.8 Y12 + 3.6 Y13 + 2.9 Y14 + 4.9 Y21 + 7.2 Y22 + 4.3 Y23 + 3.1 Y24

Subject to,

Y11 + Y12 + Y13 + Y14 = 2

Y21 + Y22 + Y23 + Y24 = 2

Y1k + Y2k = 1 for k=1,2,3,4

Yjk = {0,1}

Excel formulation:

A D E G 1 Dishwashing (3) Laundry (4) Budgeting (1) Cooking (2) Total 2 Eve -SUM(C3F3) 3 Steven |-SUM(C4F4) 4 Total E3+E4 —С3

Set Objective: $C$11 To: OValue Of: C Min Мax By Changing Variable Cells: $C$3:$F$4 Subject to the Constraints: $C$3:$F$4 bin

Solution:

Budgeting
(1)
Cooking
(2)
Dishwashing
(3)
Laundry
(4)
Total
Eve 1 0 1 0 2
Steven 0 1 0 1 2
Total 1 1 1 1
Budgeting
(1)
Cooking
(2)
Dishwashing
(3)
Laundry
(4)
Eve 4.5 7.8 3.6 2.9
Steven 4.9 7.2 4.3 3.1
Total time 18.4